12 horses are in a race;
a) what is the probability that three will win?
b) what is the probability that three will be awarded 1st, 2nd, and 3rd place?
Answer #1

A)what is the probability that three will win?

Prob that three will be win is = 3/12

0.25

b) what is the probability that three will be awarded 1st, 2nd, and 3rd place?

therfor first three winner I know but only awarded to them 1st ,2nd or 3rd

means prob is first awarded is 1/3

second awarded prob is 1/2

third is awarded prob 1/1

=1/3*1/2*1/1

=1/6

=0.17

A)0.25

B) 0.17